# CoT Referring: Improving Referring Expression Tasks with Grounded Reasoning

### v1

@@ Abstract
**footnotetext: This work was done during an internship at Adobe Research.footnotetext: Corresponding author.
@@ 1 Introduction
The advent of Multimodal Large Language Models (MLLMs) has revolutionized vision-and-language research, particularly in Referring Expression (RE) Tasks, such as Referring Expression Comprehension (REC)  (Mao et al., 2016) and Referring Expression Segmentation (RES) (also known as Referring Image Segmentation)  (Hu et al., 2016). This task is crucial as it demands a model’s ability to integrate visual and textual information seamlessly. Recent advancements, such as  (Li et al., 2024b; Rasheed et al., 2024; Ma et al., 2024; Zhang et al., 2024; Ren et al., 2024; Peng et al., 2023; Ma et al., 2023; Wang, 2022), have empowered Large Language Models (LLMs) to perform localization of specific objects in an image based on language queries.
@@ 2.1 Phrase Grounding
Phrase grounding, also known as visual grounding, aims to “ground” (i.e., localize) every relevant object in a text query in an image (Karpathy & Fei-Fei, 2015; Rohrbach et al., 2016; Plummer et al., 2016). Early works relied on merging pre-trained box proposal networks or object detectors to ground objects (Chen et al., 2020a; Li et al., 2020; Kamath et al., 2021). Some methods focus on weakly- and self-supervised techniques to reduce dependence on expensive region annotations (Arbelle et al., 2021; Shaharabany et al., 2022; He et al., 2023), while others introduce contrastive learning approaches that improve phrase–object alignment and phrase grounding capabilities (Gupta et al., 2020; Li et al., 2021; Radford et al., 2021). Phrase grounding datasets are mainly derived from caption corpora that match nouns to regions, whereas referring-expression datasets provide descriptions of a single target object.
@@ 2.3 Models for Referring Expression Tasks
Existing RE methods can be grouped into two families: specialist visual grounding models and MLLM-based models. Numerous specialist models are designed specifically for visual grounding tasks (Liu et al., 2023; Kamath et al., 2021; Chng et al., 2024; Zou et al., 2023; Wang et al., 2024; Chen et al., 2024b). These models typically employ dedicated architectures optimized for detection and segmentation.
MLLMs have seen rapid advances on vision–language tasks with detection and segmentation abilities. Recent developments (Ma et al., 2024; Ren et al., 2024; Rasheed et al., 2024; Chen et al., 2024c; Peng et al., 2023; Ma et al., 2023; Wang et al., 2023) focus on enhancing localization capabilities. Many approaches leverage instruction-following frameworks, often framing RE as instruction following with specialized heads for detection/segmentation. Some introduce special tokens (e.g., [SEG]) and require additional fine-tuning and alignment during training to handle box/mask generation. MLLMs (Ma et al., 2024; Ren et al., 2024; Rasheed et al., 2024; Chen et al., 2024c) also make use of pre-trained visual models and decoders to generate masks, such as ViT (Dosovitskiy et al., 2020) and SAM (Kirillov et al., 2023).
@@ 3.1 Data Generation Pipeline
We address these questions with a four-stage data generation pipeline, depicted in Figure 3. This pipeline leverages a sequence of large language and vision-language models (Qwen3-235B-A22B (Qwen Team, 2025) $\rightarrow$ DeepSeek-V3.1 (DeepSeek-AI Team, 2024) $\rightarrow$ Qwen2.5-VL-72B (Bai et al., 2025) $\rightarrow$ GPT-4o (Hurst et al., 2024)). To quantify complexity of a referring expression, we define the hop level of a anchor noun as its dependency distance to the target, and $L_{\max}$, the maximum hop level among all anchors, as the complexity of the expression.
@@ 3.1.1 Data Format and Representation
Notation and measures. We use [LOC] as placeholders for spatial grounding, which are later mapped to bounding boxes. The hop level (H.L.) of a noun is the shortest dependency distance to the target in the parsed graph (the target has H.L.=1). We define complexity as $L_{\max}$, the maximum hop level across nouns. Larger $L_{\max}$ indicates longer reasoning chains and more challenging compositional grounding (see Fig. 2). Following prior MLLMs (Rasheed et al., 2024; Peng et al., 2023; Ma et al., 2023; Zhang et al., 2024). Thus, we adopt an instruction-following format. A standard query like "Where is the cat on the chair?" might yield "The cat on the chair is [LOC]", whereas our CoTR format rewrites it to expose intermediate steps: "First locate the chair [LOC], then the cat on top of it [LOC]." This format promotes (1) a canonical reasoning order by decomposing the query into clear, interpretable steps and (2) gradual visual grounding by mapping each [LOC] placeholder to a bounding box, providing a strong supervisory signal for each step.
@@ 3.1.3 Stage B: Visual Grounding
Having validated the text, we proceed to ground each noun and verify the results. B.1 — Grounding. Qwen2.5-VL-72B (Bai et al., 2025) is prompted to predict one bounding box for each [LOC] placeholder. To improve grounding quality, including cases with multiple instances of the same noun (e.g., “a giraffe with its head on another giraffe”) or unclear/pronominal mentions (e.g., “theirs”, “one”), we include the relevant span of each anchor in the referring expression as context in the query prompt. This added context clarifies references and helps align textual steps with correct image regions. B.2 — Verification. We verify the generated boxes with two filters: (i) for each noun, we render its predicted box on the image and ask GPT-4o (Hurst et al., 2024) for a yes/no confirmation that the box correctly localizes the noun (we found GPT-4o to be best at visual verification but not precise localization), and (ii) for the final target object, we require $\mathrm{IoU}_{\mathrm{GT}}>0.7$ against the dataset ground truth. A sample is accepted only if both conditions are satisfied.
@@ 4.2 Model Architecture
We propose RefLM (Fig. 5), which follows a typical MLLM architecture comprising a vision encoder, a projector, and an LLM. Unlike prior models (Rasheed et al., 2024; Zhang et al., 2024; Li et al., 2024b), RefLM removes the learned mask-decoder head. Instead, it predicts visual prompts (bounding boxes and points) that provide fine-grained spatial cues, and then uses the Segment Anything Model (SAM) to generate masks. This eliminates the need to learn an additional embedding-to-mask mapping (decoding [SEG] token) between the LLM and a mask decoder, because supervision can be applied directly at the token level.
@@ 5 Experiments
Implementation Details. For RefLM-8B, we initiate our weights from LLaVA-NeXT-8B (Li et al., 2024a). For RefLM-7B, we follow the setting of OMG-LLAVA-7B (Zhang et al., 2024), using OpenCLIP’s ConvNeXT-L as the vision encoder (Liu et al., 2022), and InternLM2-7B as the LLM (Cai et al., 2024). This setup for RefLM-7B is designed to be directly comparable to other leading 7B models. For RefLM-8B, we only need to conduct instruction tuning, while for RefLM-7B, an extra alignment stage is needed before instruction tuning, as mentioned in OMG-LLAVA (Zhang et al., 2024). SAM (Kirillov et al., 2023) is employed for mask prediction. For grounding of anchors, we use bounding boxes, while for targets we utilize box-point representation to learn mask representations. Training is conducted with an initial learning rate of 1e-4 and a global batch size of 128. The perception model’s weights are frozen, and the LLM is fine-tuned using LoRA with a rank of 256 (Hu et al., 2021). All experiments are conducted on eight A100 GPUs with 40GB memory.
@@
Ferret-13B (Li, 2023) 13B 63.4 60.2 60.1 64.2 60.8 59.5 61.37
CogVLM-17B (Ding, 2023) 17B 68.3 65.7 65.5 68.8 65.2 65.2 66.85
CogCom-17B (Huang, 2023) 17B 68.1 65.5 65.3 68.7 65.0 65.1 65.29
Qwen2.5-VL-7B (Bai et al., 2025) 7B 68.1 67.2 65.6 69.2 65.8 64.5 66.97
@@ Appendix D RefCOCOm Results
SeqTR (Zhu et al., 2022) 13.9 28.2 12.1 22.8 18.1 34.7
CRIS (Wang et al., 2022) 10.6 25.4 10.1 21.2 12.9 30.0
LAVT (Yang et al., 2022) 15.3 29.9 13.2 24.4 18.7 35.5
@@
X-Decoder (Zhang et al., 2023) 16.2 29.5 13.6 23.6 20.3 33.8
SEEM (Zou et al., 2023) 16.1 29.4 13.6 23.4 20.4 33.9
UniRES (Chen et al., 2024b) 19.6 34.3 16.4 27.8 25.2 41.7

### v2

@@ 1 Introduction
The advent of Multimodal Large Language Models (MLLMs) has revolutionized vision-and-language research, particularly in Referring Expression (RE) Tasks, such as Referring Expression Comprehension (REC)  (Mao et al., 2016) and Referring Expression Segmentation (RES) (also known as Referring Image Segmentation)  (Hu et al., 2016). This task is crucial as it demands a model’s ability to integrate visual and textual information seamlessly. Recent advancements, such as  (Lai et al., 2024; Rasheed et al., 2024; Ma et al., 2024a; Zhang et al., 2024; Ren et al., 2024; Peng et al., 2023; Ma et al., 2024b; Chen et al., 2023), have empowered Large Language Models (LLMs) to perform localization of specific objects in an image based on language queries.
@@ 2.1 Phrase Grounding
Phrase grounding, also known as visual grounding, aims to “ground” (i.e., localize) every relevant object in a text query in an image (Karpathy and Fei-Fei, 2015; Rohrbach et al., 2016; Plummer et al., 2016). Early works relied on merging pre-trained box proposal networks or object detectors to ground objects (Chen et al., 2020a; Li et al., 2020; Kamath et al., 2021). Some methods focus on weakly- and self-supervised techniques to reduce dependence on expensive region annotations (Arbelle et al., 2021; Shaharabany et al., 2022; He et al., 2023), while others introduce contrastive learning approaches that improve phrase–object alignment and phrase grounding capabilities (Gupta et al., 2020; Li et al., 2021; Radford et al., 2021). Phrase grounding datasets are mainly derived from caption corpora that match nouns to regions, whereas referring-expression datasets provide descriptions of a single target object.
@@ 2.3 Models for Referring Expression Tasks
Existing RE methods can be grouped into two families: specialist visual grounding models and MLLM-based models. Numerous specialist models are designed specifically for visual grounding tasks (Liu et al., 2023b; Kamath et al., 2021; Chng et al., 2024; Zou et al., 2023b; Liu et al., 2023a; Wang et al., 2024). These models typically employ dedicated architectures optimized for detection and segmentation.
MLLMs have seen rapid advances on vision–language tasks with detection and segmentation abilities. Recent developments (Ma et al., 2024a; Ren et al., 2024; Rasheed et al., 2024; Chen et al., 2024b; Peng et al., 2023; Ma et al., 2024b; Wang et al., 2023b) focus on enhancing localization capabilities. Many approaches leverage instruction-following frameworks, often framing RE as instruction following with specialized heads for detection/segmentation. Some introduce special tokens (e.g., [SEG]) and require additional fine-tuning and alignment during training to handle box/mask generation. MLLMs (Ma et al., 2024a; Ren et al., 2024; Rasheed et al., 2024; Chen et al., 2024b) also make use of pre-trained visual models and decoders to generate masks, such as ViT (Dosovitskiy et al., 2020) and SAM (Kirillov et al., 2023).
@@ 3.1 Data Generation Pipeline
We address these questions with a four-stage data generation pipeline, depicted in Figure 3. This pipeline leverages a sequence of large language and vision-language models (Qwen3-235B-A22B (Qwen Team, 2025) $\rightarrow$ DeepSeek-V3.1 (DeepSeek-AI Team, 2024) $\rightarrow$ Qwen2.5-VL-72B (Bai and others, 2025) $\rightarrow$ GPT-4o (OpenAI, 2024)). To quantify complexity of a referring expression, we define the hop level of a anchor noun as its dependency distance to the target, and $L_{\max}$, the maximum hop level among all anchors, as the complexity of the expression.
@@ 3.1.1 Data Format and Representation
Notation and measures. We use [LOC] as placeholders for spatial grounding, which are later mapped to bounding boxes. The hop level (H.L.) of a noun is the shortest dependency distance to the target in the parsed graph (the target has H.L.=1). We define complexity as $L_{\max}$, the maximum hop level across nouns. Larger $L_{\max}$ indicates longer reasoning chains and more challenging compositional grounding (see Fig. 2). Following prior MLLMs (Rasheed et al., 2024; Peng et al., 2023; Ma et al., 2024b; Zhang et al., 2024). Thus, we adopt an instruction-following format. A standard query like "Where is the cat on the chair?" might yield "The cat on the chair is [LOC]", whereas our CoTR format rewrites it to expose intermediate steps: "First locate the chair [LOC], then the cat on top of it [LOC]." This format promotes (1) a canonical reasoning order by decomposing the query into clear, interpretable steps and (2) gradual visual grounding by mapping each [LOC] placeholder to a bounding box, providing a strong supervisory signal for each step.
@@ 3.1.3 Stage B: Visual Grounding
Having validated the text, we proceed to ground each noun and verify the results. B.1 — Grounding. Qwen2.5-VL-72B (Bai and others, 2025) is prompted to predict one bounding box for each [LOC] placeholder. To improve grounding quality, including cases with multiple instances of the same noun (e.g., “a giraffe with its head on another giraffe”) or unclear/pronominal mentions (e.g., “theirs”, “one”), we include the relevant span of each anchor in the referring expression as context in the query prompt. This added context clarifies references and helps align textual steps with correct image regions. B.2 — Verification. We verify the generated boxes with two filters: (i) for each noun, we render its predicted box on the image and ask GPT-4o (OpenAI, 2024) for a yes/no confirmation that the box correctly localizes the noun (we found GPT-4o to be best at visual verification but not precise localization), and (ii) for the final target object, we require $\mathrm{IoU}_{\mathrm{GT}}>0.7$ against the dataset ground truth. A sample is accepted only if both conditions are satisfied.
@@ 4.2 Model Architecture
We propose RefLM (Fig. 5), which follows a typical MLLM architecture comprising a vision encoder, a projector, and an LLM. Unlike prior models (Rasheed et al., 2024; Zhang et al., 2024; Lai et al., 2024), RefLM removes the learned mask-decoder head. Instead, it predicts visual prompts (bounding boxes and points) that provide fine-grained spatial cues, and then uses the Segment Anything Model (SAM) to generate masks. This eliminates the need to learn an additional embedding-to-mask mapping (decoding [SEG] token) between the LLM and a mask decoder, because supervision can be applied directly at the token level.
@@ 5 Experiments
Implementation Details. For RefLM-8B, we initiate our weights from LLaVA-NeXT-8B (Li et al., 2024). For RefLM-7B, we follow the setting of OMG-LLAVA-7B (Zhang et al., 2024), using OpenCLIP’s ConvNeXT-L as the vision encoder (Liu et al., 2022), and InternLM2-7B as the LLM (Cai et al., 2024). This setup for RefLM-7B is designed to be directly comparable to other leading 7B models. For RefLM-8B, we only need to conduct instruction tuning, while for RefLM-7B, an extra alignment stage is needed before instruction tuning, as mentioned in OMG-LLAVA (Zhang et al., 2024). SAM (Kirillov et al., 2023) is employed for mask prediction. For grounding of anchors, we use bounding boxes, while for targets we utilize box-point representation to learn mask representations. Training is conducted with an initial learning rate of 1e-4 and a global batch size of 128. The perception model’s weights are frozen, and the LLM is fine-tuned using LoRA with a rank of 256 (Hu et al., 2021). All experiments are conducted on eight A100 GPUs with 40GB memory.
@@
Ferret-13B (You et al., 2023) 13B 63.4 60.2 60.1 64.2 60.8 59.5 61.37
CogVLM-17B (Wang et al., 2023a) 17B 68.3 65.7 65.5 68.8 65.2 65.2 66.85
CogCom-17B (Qi et al., 2024) 17B 68.1 65.5 65.3 68.7 65.0 65.1 65.29
Qwen2.5-VL-7B (Bai and others, 2025) 7B 68.1 67.2 65.6 69.2 65.8 64.5 66.97
@@ Appendix D RefCOCOm Results
SeqTR [Zhu et al., 2022] 13.9 28.2 12.1 22.8 18.1 34.7
CRIS [Wang et al., 2022] 10.6 25.4 10.1 21.2 12.9 30.0
LAVT [Yang et al., 2022] 15.3 29.9 13.2 24.4 18.7 35.5
@@
X-Decoder [Zou et al., 2023a] 16.2 29.5 13.6 23.6 20.3 33.8
SEEM [Zou et al., 2023b] 16.1 29.4 13.6 23.4 20.4 33.9
UniRES [Wang et al., 2024] 19.6 34.3 16.4 27.8 25.2 41.7